Midlands Inc. had a bad year in 2019. For the first time in its history, it operated at a loss. The companys income statement showed the following results from selling 75,000 units of product: net sales $ 1,500,000; total costs and expenses $ 1,900,000; and net loss $400,000. Costs and expenses consisted of the following. Total Variable Fixed 1,240,000 $755,000 $485,000 515,000 90,000 425,000 Administrative expenses 145,000 55,000 90,000 $1,900,000 $900,000 $1,000,000 Cost of goods sold Selling expenses Management is considering the following independent alternatives for 2020. Increase unit selling price 20% with no change in costs and expenses. Change the compensation of salespersons from fixed annual salaries totaling $195,000 to total salaries of $35,000 plus a 5% commission on net sales. Purchase new high-tech factory machinery that will change the proportion between variable and fixed cost of goods sold to 50:50 1. 2. 3.(a) Compute the break-even point in dollars for 2019. (Round contribution margin ratio to 4 decimal places e.g. 0.2512 and final answer to O decimal places, e.g. 2,510.) Break-even point $ (b) Compute the break-even point in dollars under each of the alternative courses of action for 2020. (Round contribution margin ratio to 3 decimal places e.g. 0.251 and final answers to O decimal places, e.g. 2,510.) Break-even point 1. Increase selling price 2. Change compensation$ 3. Purchase machinery $ Which course of action do you recommend?
Answer #1

Answers

  • All working forms part of requirement
  • Requirement ‘a’

Break Even point = $ 2,500,000

A

Net Sales

$         1,500,000

B

Variable Cost

$             900,000

C = A - B

Contribution margin

$             600,000

D = C/A

CM Ratio

0.4000

E

Total Fixed Cost

$         1,000,000

F = E/D

Break Even in $

$         2,500,000

  • Requirement ‘b’

Break Even point

1. Increasing selling price

$        2,000,000

2. Change compensation

$        2,400,000

3. Purchase machinery

$        2,316,327

--Which course is recommended?
Answer = #1 Increasing Selling price, because this is reducing Break Even level and cost is unchanged.

--Workings for above

#1: Increasing selling price

A

Current Sales

$         1,500,000

B = A x 20%

Increase in Sale

$             300,000

C = A + B

2020's sale

$         1,800,000

D

Variable Cost

$             900,000

E = C - D

Contribution margin

$             900,000

F = E/C

CM Ratio

0.5000

G

Total Fixed Cost

$         1,000,000

H = G/F

Break Even in $

$         2,000,000

#2 Change Compensation

Variable Cost

Fixed Cost

Current

$             900,000

$         1,000,000

Old compensation

$          (195,000)

New compensation

$               75,000

$               35,000

New Costs for 2020

$             975,000

$             840,000

A

2020's sale

$         1,500,000

B

Variable Cost

$             975,000

C = A - B

Contribution margin

$             525,000

D = C/A

CM Ratio

0.350

E

Total Fixed Cost

$             840,000

F = E/D

Break Even in $

$         2,400,000

#2 Purchase machinery

Variable Cost

Fixed Cost

Current

$      900,000

$               1,000,000

Old Cost of Goods Sold

$   (755,000)

$                 (485,000)

New Cost of Goods Sold

$      620,000

$                   620,000

New Costs for 2020

$      765,000

$               1,135,000

A

2020's sale

$               1,500,000

B

Variable Cost

$                   765,000

C = A - B

Contribution margin

$                   735,000

D = C/A

CM Ratio

0.490

E

Total Fixed Cost

$               1,135,000

F = E/D

Break Even in $

$               2,316,327